OpenStudy (cicilybailey14):

Man Wrinkle spent one-fourth of his life as a boy, one-eighth as a youth, and one-half as an active man. If Man Wrinkle spent 12 years as an old man, then how many years did he spend as an active man? @Samuel917

OpenStudy (anonymous):

80 years.

OpenStudy (anonymous):

THATS HOW DAMN GOOD I AM

OpenStudy (anonymous):

Its too bad...

OpenStudy (anonymous):

I'm so good.

OpenStudy (anonymous):

Wtf??

OpenStudy (cicilybailey14):

fr^

OpenStudy (anonymous):

Ya totally..

OpenStudy (cicilybailey14):

@Samuel917 how do I know you didn't look it up. Chill

OpenStudy (anonymous):

48 years

OpenStudy (anonymous):

Oh. That's true. I'm just really good at thees. And I like to brag.

OpenStudy (anonymous):

Woooooo medals for dayyyyss

OpenStudy (anonymous):

Samuel your dumb. how would someone be an elder for 80 years.

OpenStudy (anonymous):

Go hump a tree.

OpenStudy (cicilybailey14):

omfg^^^ I died @Tomas.Moreno

OpenStudy (cicilybailey14):

@AleshGames ur right

OpenStudy (anonymous):

80/2 = 40 years. Your answer. Sorry. I though everyone would be smart enough to understand the question. My bad.

OpenStudy (anonymous):

Fraction of life as a boy = 1/4 Fraction of life as a youth = 1/8 Fraction of life as an active man = 1/2 Fraction of life as boy, youth and active man = 1/4 + 1/8 + 1/2 = (2 + 1 + 4)/8 = 7/8 Fraction of life as an old man = 1 − 7/8 = 1/8 Thus, one-eighth of Man Wrinkle's life (as an old man) is 10 years. So, Man Wrinkle's Age = 80 years.

OpenStudy (anonymous):

Fraction of life as a boy = 1/4 Fraction of life as a youth = 1/8 Fraction of life as an active man = 1/2 Fraction of life as boy, youth and active man = 1/4 + 1/8 + 1/2 = (2 + 1 + 4)/8 = 7/8 Fraction of life as an old man = 1 − 7/8 = 1/8 Thus, one-eighth of Man Wrinkle's life (as an old man) is 12 years. So, Man Wrinkle's Age = 96 years. It may be noted that: Life as boy = 96/4 = 24 years. Life as youth = 96/8 = 12 years. Life as active man = 96/2 = 48 years. Life as old man = 96/8 = 12 years. The problem may also be solved by setting up the following equation: a/4 + a/8 + a/2 + 12 = a where a denotes Man Wrinkle's age in years. The equation may be solved as shown below. 7a/8 + 12 = a 12 = a − 7a/8 = a/8 a/8 = 12 or a = 96. Therefore, life spent as active man = 96/2 = 48 years.

OpenStudy (anonymous):

;) The correct answer.

OpenStudy (cicilybailey14):

@Samuel917 http://www.syvum.com/cgi/online/fillin.cgi/teasers/wrinkle1.tdf?0 go to this site

OpenStudy (cicilybailey14):

type in 80. it's wrong than type in 48 its right

OpenStudy (anonymous):

The problem may also be solved by setting up the following equation: a/4 + a/8 + a/2 + 11 = a where a denotes Man Wrinkle’s age in years. The equation may be solved as shown below. 7a/8 + 11 = a 11 = a − 7a/8 = a/8 a/8 = 11 or a = 88. Therefore, life spent as active man = 88/2 = 44 years.

OpenStudy (anonymous):

I win

OpenStudy (cicilybailey14):

Who says?

OpenStudy (anonymous):

Law. Order. Logic.

OpenStudy (cicilybailey14):

Your vexatious